# Tumor-anchored deep feature random forests for out-of-distribution detection in lung cancer segmentation

### v1

@@ 5.3.2 RF-Deep extracts OOD data-agnostic and specific features for improved separability
We used SHAP (SHapley Additive exPlanations) (lundberg2017unified) to identify which features drive RF-Deep and RF-Radiomics predictions, revealing how the two approaches differ in their decision-making mechanisms. Deep features (Figure 6) show broad, dataset-specific spread of important dimensions, with some features reused across multiple datasets (Ft 65, Ft 131, Ft 179, Ft 319), suggesting they capture general distributional properties relevant to distinguishing ID from OOD scans. Other features activate selectively for specific distribution shifts (e.g., Ft 5 and Ft 1381 for far-OOD), indicating that RF-Deep adapts its decision boundaries to the characteristics of each OOD scenario rather than relying on a fixed, universal subset of features. Notably, only a small fraction of the 1,488 encoder features appear consistently useful across OOD datasets, underscoring the need for outlier data exposure, especially in near-OOD scenarios where subtle feature combinations are required to distinguish concept shifts.
@@ 5.3.3 Why logit-based spatial uncertainties are less accurate than RF-Deep?
MaxLogit scores of ID and OOD scans show strong overall separation (Mann-Whitney U tests, all p << 0.001, effect sizes ranging from r = -0.75 for MIDRC C19 to r = -0.90 for KiTS) due to higher overall MaxLogit scores for ID cases than individual OOD datasets (Table 2). However, the scores have substantially overlapping ranges with ID (ID: 0.12–4.33, RSNA PE: 0.09–3.67, MIDRC C19: 0.30–2.70).
@@
Figure 8 provides a visualization of the MaxLogit spatial uncertainties for four representative cases. The density-normalized histograms depict the distribution of MaxLogit values in the interior and boundary regions of the detected lesions. The interior and boundary regions show very good separation for the ID case (row 1) and sufficiently different spatial distribution for the second case (row 2), enabling correct detection of OOD for this case. However, the third case (row 3) presents a challenge wherein the spatial uncertainty distribution closely resembles that of the ID case, resulting in incorrect classification. In contrast, RF-Deep correctly identified both OOD scans, demonstrating its superior ability to capture contextual features beyond local spatial patterns. Of note, both RF-Deep and MaxLogit-based OOD prediction was wrong for the fourth case (row 4), presenting a limiting case for either method.
An additional limitation of logit-based methods is their inherent scale invariance: MaxLogit scores reflect local appearance characteristics independent of lesion size as shown by the distributions for lesions of varying sizes in Figure 8. This is because MaxLogit computation treats voxels identically based on softmax probabilities, ignoring volumetric context that could differentiate true tumors from anatomically implausible detections. This scale invariance poses a critical bottleneck for OOD detection, as many false-positive detections on OOD scans manifest as small, scattered regions with locally high confidence but globally inconsistent spatial distributions. A method like RF-Deep on the other hand, aggregates features from regions encompassing detections, reducing the impact of tumor size variations.
@@ 5.4.1 Effect of classifier selection
Table 3 compares different classifiers applied to the same set of encoder features. While linear probing and multilayer perceptron (MLP) classifiers, both implemented with default settings from scikit-learn, provide reasonable separability on the far-OOD cohorts, their performance degrades notably on the near-OOD cohorts (AUROC $\approx 92--93\%$, FPR95 $\approx 28--30\%$). In contrast, random forests achieve the best performance across all cohorts, highlighting the advantage of tree-based non-linear classifiers for capturing complex decision boundaries, particularly when distribution shifts are subtle (Figure 4(a)). The perfect performance on KiTS and PancreasCT (AUROC = 100%, FPR95 $\approx$ 0%) reflects the very easy nature of far-OOD detection when anatomical regions differ substantially from the training distribution.
@@ 5.4.2 Generalization to unseen OOD categories
We evaluated RF-Deep’s ability to generalize to previously unseen OOD categories using four additional training strategies (Table 4): (i) dataset-specific (DS; default), (ii) ensemble, (iii) unified, (iv) leave-one-dataset-out (LODO), and LODO+. DS trained a separate RF classifier for each OOD dataset, representing an upper-bound setting where the OOD type was known during development and evaluation. Ensemble trained the same four dataset-specific classifiers but averaged their predictions at test time, assuming no knowledge of the test domain during testing. Unified replaced the per-dataset training with a single RF classifier trained on ID versus all OOD datasets combined. LODO followed the standard OOD paradigm by training on ID plus three OOD datasets and evaluating on the held-out fourth, directly measuring generalization to novel OODs. Finally, LODO+ extended LODO by incorporating sub-sampled, non-tumor-centric ROIs to increase anatomical diversity during training.
@@ 5.5 Design choices and sensitivity analysis
To understand the design choices of our proposed OOD detector, we conducted a series of controlled ablations on the number of ROIs, their crop sizes, the extent of labeled (supervised) data, and the encoder stages used for feature extraction. Figure 9(a) shows that OOD detection benefits from richer local context around predicted tumor regions, but only up to a point. With $n=1$, performance is unstable, reflected in wider confidence intervals and elevated FPR95, suggesting that a single tumor-anchored crop may miss informative variation in tumor surroundings. Increasing the number of ROIs from one to four improved performance on near-OOD cases, indicating that additional samples may better disambiguate similarly appearing tumor (ID) from non-cancerous structures. Beyond four ROIs, gains plateau and in some cases regress (e.g., MIDRC C19), implying diminishing returns and potential noise injection from redundant or low-value ROIs. Far-OOD datasets did not exhibit performance gains with increasing number of ROIs. Therefore, we used $n=4$ to capture meaningful contextual diversity while maintaining computational efficiency.
@@
Figure 9(b) analyzes the effect of tumor-anchored 3D crop size on OOD detection performance. Performance remained relatively stable on the far-OOD cohorts across crop sizes, with slight preference for $128\times 128\times 128$. On the near-OOD cohorts, both smaller (<128<128) and larger (>128>128) ROIs led to performance degradation, suggesting that the optimal contextual window balances tumor-centered information with surrounding anatomical context.
@@
Figure 9(c) shows that RF-Deep’s OOD detection performance improved with the number of labeled examples, using equal numbers of ID and OOD scans (5/5, 10/10, up to 20/20, corresponding to less than 30% of available data), after which performance plateaued. With only 20 ID and 20 OOD labeled scans, RF-Deep achieved AUROC >> 90 and FPR95 << 30% on average across all OOD datasets. Performance gains were more noticeable on the near-OOD cohorts than the far-OOD cohorts, indicating that detection on subtle shifts in the same anatomical region is trickier and requires more examples than larger shifts across different anatomical regions. Variance decreases consistently as more samples are added, with the strongest variance reduction observed on near-OOD cohorts.
@@
Figure 9(d) shows individual Swin Transformer stage contributions. Far-OOD cohorts were relatively insensitive to individual stage selection (achieving AUROC >> 98 even when using features from a single stage), whereas the near-OOD counterparts showed strong sensitivity to early stages, with Stage 1 and Stage 2 features being particularly important (AUROC $\approx$ 88–90 individually). Concatenating features across all stages yielded the best performance, confirming that hierarchical representations are essential for robust OOD detection. This finding aligns with the SHAP analysis (Section 5.3.2), which showed that RF-Deep adaptively leverages features from different abstraction levels depending on the OOD scenario.
6 Discussion
We introduced a lightweight approach for instance-level OOD detection for lung tumor segmentation involving near-OOD and far-OOD cases. Our approach, termed RF-Deep, leveraged the advantage of SSL pretrained transformer encoders to extract image aquisition robust features, while using RF’s feature selection strategy to extract representations that distinguished subtle differences in pathologies occurring in the same disease site as the task model. Our approach improved performance over multiple current OOD detection methods as well as handcrafted radiomics-based RF classifiers under all evaluated datasets. We also found that RF-Deep approach was agnostic to the choice of pretraining approach as well as model size, thus providing a plug-and-play approach for leveraging an existing task-specific DL method for OOD detection.
@@ 6 Discussion
Analysis of why RF-Deep outperformed other methods showed improved separability of ID from OOD datasets using deep compared to handcrafted radiomic features. Importantly, despite using the same deep features, Mahalanobis distance-based approach was less accurate than RF-Deep, indicating the need for outlier exposure and feature selection performed with random forests to extract a robust non-linear decision boundary. RF-Deep improved over logits-based methods, indicating that over-reliance on the ID model to extract confidence boundaries may be unreliable in OOD scans. Although this issue could be addressed to an extent through confidence calibration, prior studies show limitations of such methods (Larrazabal2023MEEP; mehrtash2020confidence).
Our approach has a few limitations. First, RF-Deep performs OOD detection only in scans where a segmentation was generated. Reduction of missing segmentations on ID scans is an altogether different problem and beyond scope of this work. Second, our approach was applied to binary tumor segmentation tasks. Extension to multi-class segmentation, with richer datasets, is straightforward and left for future work. Finally, the evaluated OOD datasets captured some of the key failure modes, but covering all possible pathologies is impractical due to lack of such datasets. Methodologies including continual learning with aggressive data augmentation for incorporating newly discovered failure modes are the next area of development.
7 Conclusion
@@
We introduced RF-Deep, a post-hoc out-of-distribution (OOD) detection framework for segmentation models that leverages hierarchical encoder features and random forest classifiers to identify concept drifts. Across near-OOD and far-OOD cohorts, RF-Deep consistently outperformed logit-based methods and radiomics feature-based classifiers. The method is practical as it seamlessly integrates into existing segmentation pipelines without any modification, works across different pretraining strategies and architectures, and provides interpretable predictions through feature attribution.
Acknowledgment
This research was partially supported by the NCI R01CA258821 and the Memorial Sloan Kettering Cancer Center Support Grant/Core Grant NCI P30CA008748.
@@ Appendix A Pretraining strategies descriptions
SMIT (jiang2022self) consists of masked voxel token prediction, masked patch token distillation, and image token distillation. Masked voxel token prediction is a dense regression task of the intensities within masked patches and is optimized by minimizing the $\ell_{1}$ norm between the predicted and ground-truth patches, akin to SimMIM (xie2022simmim). Self-distillation (antti2017meanteacher) was required for the next two tasks and was performed using an exponentially moving average (EMA) teacher model with identical architecture as the student, similar to iBOT (zhou2021ibot). Masked patch token distillation aligns the voxel token representations learned by the student with those extracted by the teacher from unmasked sequences, optimized with a temperature-scaled cross-entropy loss. For image token distillation, the cross-entropy between global volume embeddings of the student and teacher was minimized. After pretraining, the teacher was discarded and the student encoder was retained for fine-tuning.
Similarly, Swin UNETR (tang2022self) employs three pretraining tasks. The primary pretext task is masked volume reconstruction, where a proportion of 3D input patches are randomly masked and the model is trained to regress their voxel intensities from visible context. Unlike SimMIM, which leverages positional encodings to aid masked token prediction, this reconstruction task provided no explicit positional information, forcing the model to learn semantically plausible inpainting from local context. 3D rotation prediction enforces invariance by classifying the degree of rotation applied along the $z$-axis. Finally, contrastive coding encouraged representation learning by maximizing cosine similarity between augmented views of the same volume while minimizing similarity for different volumes. Unlike SMIT, which requires an EMA teacher, Swin UNETR used only a student model. The pretrained encoder was directly retained for fine-tuning.
@@ Appendix B Logit-based approaches formulae
The tumor segmentation problem can be broken down into a per-voxel binary classification problem, in which we mean-aggregate the scores of the positive class to summarize the scan. In-distribution (ID) scans tended to yield more confident outputs, whereas out-of-distribution (OOD) scans produced less confident or spurious predictions. Following the convention of treating OOD as the positive class (hendrycks17baseline; choi2023conservative), we negated each score so that higher values correspond to greater likelihood of being OOD. Let $f^{i}(x)$ denote the logit for class $i\in\{0,1\}$ at voxel $x$, and define the softmax score as
@@
  • 1.
    MaxSoftmax (hendrycks17baseline):

### v2

@@ 5.3.2 RF-Deep extracts OOD data-agnostic and specific features for improved separability
We used SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017) to identify which features drive RF-Deep and RF-Radiomics predictions, revealing how the two approaches differ in their decision-making mechanisms. Deep features (Figure 5) show broad, dataset-specific spread of important dimensions, with some features reused across multiple datasets (Ft 65, Ft 131, Ft 179, Ft 319), suggesting they capture general distributional properties relevant to distinguishing ID from OOD scans. Other features activate selectively for specific distribution shifts (e.g., Ft 5 and Ft 1381 for far-OOD), indicating that RF-Deep adapts its decision boundaries to the characteristics of each OOD scenario rather than relying on a fixed, universal subset of features. Notably, only a small fraction of the 1,488 encoder features appear consistently useful across OOD datasets, underscoring the need for outlier exposure, especially in near-OOD scenarios where subtle feature combinations are required to distinguish concept shifts.
@@ 5.3.3 Why logit-based spatial uncertainties are less accurate than RF-Deep?
MaxLogit scores of ID and OOD scans show strong overall separation (Mann-Whitney U tests, all p << 0.001, effect sizes ranging from r = -0.75 for MIDRC C19 to r = -0.90 for KiTS) due to higher overall MaxLogit scores for ID scans than individual OOD datasets (Table 2). However, the scores have substantially overlapping ranges with ID (ID: 0.12–4.33, RSNA PE: 0.09–3.67, MIDRC C19: 0.30–2.70).
@@
Figure 7 provides a visualization of the MaxLogit spatial uncertainties for four representative scans. The density-normalized histograms depict the distribution of MaxLogit values in the interior and boundary regions of the detected lesions. The interior and boundary regions show very good separation for the ID case (row 1) and sufficiently different spatial distribution for the second case (row 2), enabling correct detection of OOD for this case. However, the third case (row 3) presents a challenge wherein the spatial uncertainty distribution closely resembles that of the ID case, resulting in incorrect classification. In contrast, RF-Deep correctly identified both OOD scans, demonstrating its superior ability to capture contextual features beyond local spatial patterns. Of note, both RF-Deep and MaxLogit-based OOD prediction was wrong for the fourth case (row 4), presenting a limiting case for either method.
An additional limitation of logit-based methods is their inherent scale invariance: MaxLogit scores reflect local appearance characteristics independent of lesion size as shown by the distributions for lesions of varying sizes in Figure 7. This is because MaxLogit computation treats voxels identically based on their pre-softmax logit values, ignoring volumetric context that could differentiate true tumors from anatomically implausible detections. This scale invariance poses a critical bottleneck for OOD detection, as many false-positive detections on OOD scans manifest as small, scattered regions with locally high confidence but globally inconsistent spatial distributions. RF-Deep, on the other hand, aggregates features from regions encompassing detections, reducing the impact of tumor size variations.
@@ 5.4.1 Effect of classifier selection
Table 3 compares different classifiers applied to the same set of encoder features. While linear probing and multilayer perceptron (MLP) classifiers, both implemented with default settings from scikit-learn, provide reasonable separability on the far-OOD datasets, their performance degrades notably on the near-OOD datasets (AUROC $\approx 92--93\%$, FPR95 $\approx 28--30\%$). In contrast, random forests achieve the best performance across all datasets, highlighting the advantage of tree-based non-linear classifiers for capturing complex decision boundaries, particularly when distribution shifts are subtle (Figure 4(a)). The perfect performance on KiTS and PancreasCT (AUROC = 100%, FPR95 $\approx$ 0%) reflects the very easy nature of far-OOD detection when anatomical regions differ substantially from the training distribution.
@@ 5.4.2 Generalization to unseen OOD categories
We evaluated RF-Deep’s ability to generalize to previously unseen OOD categories using four additional training strategies (Table 4): (i) ensemble, (ii) unified, (iii) leave-one-dataset-out (LODO), and (iv) LODO+. Dataset-specific (DS; baseline) trained a separate RF classifier for each OOD dataset, representing an upper-bound setting where the OOD type was known during development and evaluation. Ensemble trained the same four dataset-specific classifiers but averaged their predictions at test time, assuming no knowledge of the test domain during testing. Unified replaced the per-dataset training with a single RF classifier trained on ID versus all OOD datasets combined. LODO followed the standard OOD paradigm by training on ID plus three OOD datasets and evaluating on the held-out fourth, directly measuring generalization to novel OODs. Finally, LODO+ extended LODO by incorporating randomly sampled non-tumor-centric ROIs to increase anatomical diversity during training.
@@ 5.5 Design choices and sensitivity analysis
To understand the design choices of our proposed OOD detector, we conducted a series of controlled ablations on the number of ROIs, their crop sizes, the extent of labeled (supervised) data, and the encoder stages used for feature extraction. Figure 8(a) shows that OOD detection benefits from richer local context around predicted tumor regions, but only up to a point. With $n=1$, performance is unstable, reflected in wider confidence intervals and elevated FPR95, suggesting that a single tumor-anchored crop may miss informative variation in tumor surroundings. Increasing the number of ROIs from one to four improved performance on near-OOD scans, indicating that additional samples may better disambiguate similarly appearing tumor (ID) from non-cancerous structures. Beyond four ROIs, gains plateau or occasionally regress (e.g., MIDRC C19), likely due to noise injection from ROIs sampling redundant or low-value crops. Far-OOD datasets did not exhibit performance gains with increasing number of ROIs. Therefore, we used $n=4$ to capture meaningful contextual diversity while maintaining computational efficiency.
@@
Figure 8(b) analyzes the effect of tumor-anchored 3D crop size on OOD detection performance. Performance remained relatively stable on the far-OOD datasets across crop sizes, with slight preference for $128\times 128\times 128$. On the near-OOD datasets, both smaller (<128<128) and larger (>128>128) ROIs led to performance degradation, suggesting that the optimal contextual window balances tumor-centered information with surrounding anatomical context.
@@
Figure 8(c) shows that RF-Deep’s OOD detection performance improved with the number of labeled examples, using equal numbers of ID and OOD scans (5/5, 10/10, up to 20/20, corresponding to less than 30% of available data), after which performance plateaued. With only 20 ID and 20 OOD labeled scans, RF-Deep achieved AUROC >> 90 and FPR95 << 30% on average across all OOD datasets. Performance gains were more noticeable on the near-OOD datasets than the far-OOD datasets, indicating that detection on subtle shifts in the same anatomical region is trickier and requires more examples than larger shifts across different anatomical regions. Variance decreases consistently as more samples are added, with the strongest variance reduction observed on near-OOD datasets.
@@
Figure 8(d) shows individual Swin Transformer stage contributions. Far-OOD datasets were relatively insensitive to individual stage selection (achieving AUROC >> 98 even when using features from a single stage), whereas the near-OOD counterparts showed strong sensitivity to early stages, with Stage 1 and Stage 2 features being particularly important (AUROC $\approx$ 88–90 individually). Concatenating features across all stages yielded the best performance, confirming that hierarchical representations are essential for robust OOD detection. This finding aligns with the SHAP analysis (Section 5.3.2), which showed that RF-Deep adaptively leverages features from different abstraction levels depending on the OOD scenario.
6 Discussion and conclusion
@@
The main finding of this work is that effective near-OOD detection requires both discriminative features and learned decision boundaries. Despite using identical deep features, the unsupervised Mahalanobis distance (no OOD scans) performed substantially worse than RF-Deep on the near-OOD datasets, demonstrating that RF-Deep’s outlier exposure adoption and non-linear feature selection are essential for distinguishing pathologies within the same anatomical site. RF-Deep improved over logit-based methods, indicating that over-reliance on the ID model to extract confidence boundaries may be unreliable in OOD scans. Although this issue could be addressed to an extent through confidence calibration, prior studies show limitations of such methods (Larrazabal et al., 2023; Mehrtash et al., 2020). RF-Deep overcomes this limitation by combining SSL-pretrained backbones for segmentation with lightweight random forest classifiers, an approach that is agnostic to the choice of pretraining approach as well as model size, thus providing a plug-and-play approach for leveraging an existing task-specific DL method for OOD detection.
Potential biases and mitigation. To mitigate data-specific biases, we employed: (a) SSL pretraining on 10,432 diverse CT scans encompassing multiple institutions and scanner manufacturers, (b) separate training and testing datasets from different patient datasets providing external validation, and (c) disaggregated evaluation across scanner manufacturer, reconstruction kernels, and contrast agent presence (Figure 3(b)).
Limitations. RF-Deep depends on the segmentation model producing predictions; for the 10% of OOD scans with no foreground output, alternative evaluation strategies (e.g., flagging the absent predictions) would be needed as the datasets expand. While LODO experiments demonstrate generalization to unseen OOD categories (Table 4), near-OOD performance degrades when the held-out category differs substantially from training outliers. Finally, our evaluation was limited to the task of single-label lung cancer segmentation, and generalization to other anatomical sites and multi-label tumors needs to be evaluated.
@@
In summary, we introduced RF-Deep, a post-hoc out-of-distribution (OOD) detection framework for segmentation models that leverages hierarchical encoder features and random forest classifiers to identify concept drifts. Across near-OOD and far-OOD datasets, RF-Deep consistently outperformed logit-based methods and radiomics feature-based classifiers. Future work will explore extension to multi-label segmentations and integration of the absent-prediction signal evaluation for improved robustness.
Broader Impact Statement
This work addresses a critical gap between research-grade AI performance and clinical deployment requirements for medical image segmentation. By providing reliable OOD detection, RF-Deep can help prevent confidently incorrect predictions from reaching clinical workflows, potentially reducing harm to patients from automated systems that fail silently on out-of-distribution inputs. However, users should be aware that no OOD detection method is perfect; the evaluated datasets capture key failure modes, but covering all possible pathologies is impractical. Clinical deployment should include human oversight and validation on institution-specific data distributions before relying on model predictions for patient care decisions.
@@ Appendix A Pretraining strategies descriptions
SMIT (Jiang et al., 2022) consists of masked voxel token prediction, masked patch token distillation, and image token distillation. Masked voxel token prediction is a dense regression task of the intensities within masked patches and is optimized by minimizing the $\ell_{1}$ norm between the predicted and ground-truth patches, akin to SimMIM (Xie et al., 2022). Self-distillation (Tarvainen and Valpola, 2017) was required for the next two tasks and was performed using an exponentially moving average (EMA) teacher model with identical architecture as the student, similar to iBOT (Zhou et al., 2022). Masked patch token distillation aligns the voxel token representations learned by the student with those extracted by the teacher from unmasked sequences, optimized with a temperature-scaled cross-entropy loss. For image token distillation, the cross-entropy between global volume embeddings of the student and teacher was minimized. After pretraining, the teacher was discarded and the student encoder was retained for fine-tuning.
Similarly, Swin UNETR (Tang et al., 2022) employs three pretraining tasks. The primary pretext task is masked volume reconstruction, where a proportion of 3D input patches are randomly masked and the model is trained to regress their voxel intensities from visible context. Unlike SimMIM, which leverages positional encodings to aid masked token prediction, this reconstruction task provided no explicit positional information, forcing the model to learn semantically plausible inpainting from local context. 3D rotation prediction enforces invariance by classifying the degree of rotation applied along the $z$-axis. Finally, contrastive coding encouraged representation learning by maximizing cosine similarity between augmented views of the same volume while minimizing similarity for different volumes. Unlike SMIT, which requires an EMA teacher, Swin UNETR used only a student model. The pretrained encoder was directly retained for fine-tuning.
@@ Appendix B Logit-based approaches formulae
The tumor segmentation problem can be broken down into a per-voxel binary classification problem, in which we mean-aggregate the scores of the positive class to summarize the scan. In-distribution (ID) scans tended to yield more confident outputs, whereas out-of-distribution (OOD) scans produced less confident or spurious predictions. Following the convention of treating OOD as the positive class (Hendrycks and Gimpel, 2017; Choi et al., 2024), we negated each score so that higher values correspond to greater likelihood of being OOD. Let $f^{i}(x)$ denote the logit for class $i\in\{0,1\}$ at voxel $x$, and define the softmax score as
@@
  • MaxSoftmax (Hendrycks and Gimpel, 2017):